# Rethinking Independent Cross-Entropy Loss For Graph-Structured Data

### v1

@@ 1 Introduction
The joint distribution disperses prediction densities of a node over a larger label space, thereby relieving the over-confident decision making. We comprehensively test on small, large, class-imbalanced, and heterophilic graphs. The experiments on 12 datasets and 7 backbone models consistently validate the substantial generalization capability of joint-cluster distribution learning. The code is available at: https://github.com/MR9812/Joint-Cluster-Supervised-Learning
@@ 3 Joint-cluster Supervised Learning
denotes the continuous distribution of cluster representation. In practice, since the node is nearly independent to the other clusters, the approximation deduction in Eq.(6) only uses the dwelling cluster feature to obtain the marginalized distribution. As illustrated in Figure 1, given the two-dimensional prediction corresponding to truth , we sum the prediction scores row wisely to estimate . Unlike the standard supervised learning, during model inference, we make use of the cluster reference signal to reason the node classes rationally and robustly. This merit is functionally similar to the in-context learning explored recently (Min et al., 2021), where a set of data-label pairs are concatenated with input to guide the language model to make more accurate decisions. As empirically studies in Appendix H, in graph data, we observe the joint distribution modeling provides better node classification accuracy compared with the simple concatenation. We provide the more detailed explanations about joint-cluster supervised learning in Appendix E.
@@ 4 Related Work
In the realm of GNNs, label propagation works on the assumption that nodes connected by an edge are likely to share the same label (Wang & Leskovec, 2020; Zhou et al., 2023). It propagates node labels along with edge weights throughout the graph (Wang & Leskovec, 2021; Xie et al., 2022b), and then infer the unlabeled nodes effectively. Difference compared to existing work: While the label propagation often infer nodes without considering the node features (e.g., at post-processing phase), our joint-cluster loss could work with any GNN backbones to comprehensively learn the structure, feature, and label information end-to-end.
@@
To leverage the label correlation, there has been previous work in combining conditional random fields (CRF) with GNNs. CGNF (Ma et al., 2018) learns the pairwise label correlation with pairwise energy function, which is optimized to train GNNs. CRF-GNNs inserts CRF layer between the graph convolutional layers, which regularizes GNNs to preserve the label dependencies among nodes (Gao et al., 2019). SPN (Qu et al., 2022) models the local label correlation of each linked node pair via nodeGNN and edgeGNN and takes the edges in graph as input to propagate all the pairwise label correlations along edges. Difference compared to existing work: ① While CRF-based methods focus on modeling the local label correlation of every linked node pair, we learn the global joint distribution of node and its cluster. ② Our proposals show promising training scalability and inference efficiency. The CRF-based methods take the whole graph as input to propagate all the pairwise label correlations along edges. In contrast, we train and infer the joint distribution of target node only with one reference signal (i.e., cluster), which allows the batch training on large graphs (e.g., Amazon with millions of nodes).
@@ 5 Experiments
In this section, we evaluate the proposed joint-cluster learning framework on 12 public datasets over 7 backbone models, to validate the effectiveness and robustness.
@@ Appendix A Algorithm
The detailed description of our proposed joint-cluster learning framework provided.

### v2

@@ 1 Introduction
The joint distribution disperses prediction densities of a node over a larger label space, thereby relieving the over-confident decision making. We comprehensively test on small, large, class-imbalanced, and heterophilic graphs. The experiments on 12 datasets and 7 backbone models consistently validate the substantial generalization capability of joint-cluster distribution learning.
@@ 3 Joint-cluster Supervised Learning
denotes the continuous distribution of cluster representation. In practice, since the node is nearly independent to the other clusters, the approximation deduction in Eq.(6) only uses the dwelling cluster feature to obtain the marginalized distribution. As illustrated in Figure 1, given the two-dimensional prediction corresponding to truth , we sum the prediction scores row wisely to estimate . Unlike the standard supervised learning, during model inference, we make use of the cluster reference signal to reason the node classes rationally and robustly. This merit is functionally similar to the in-context learning strategy explored recently (Min et al., 2021), where a set of data-label pairs are concatenated with input to guide the language model to make more accurate decisions. As empirically studies in Appendix H, in graph-structured data, we observe the joint distribution modeling provides better node classification accuracy compared with the simple concatenation. Furthermore, we provide the more detailed explanations about joint-cluster supervised learning in Appendix E.
@@ 4 Related Work
In the realm of GNNs, label propagation works on the assumption that nodes connected by an edge are likely to share the same label (Shi et al., 2020; Wang & Leskovec, 2020; Zhou et al., 2023). It propagates node labels along with edge weights throughout the graph (Wang & Leskovec, 2021; Xie et al., 2022b), and then infer the unlabeled nodes effectively. Difference compared to existing work: While the label propagation often infer nodes without considering the node features (e.g., at post-processing phase), our joint-cluster learning framework could work with any GNN backbones to comprehensively learn the structure, feature, and label information end-to-end.
@@
To leverage the label correlation in node classification, there has been previous work in combining conditional random fields (CRF) with GNNs. CGNF (Ma et al., 2018) learns the pairwise label correlation with pairwise energy function, a specific expression form of CRF, which is optimized to train GNNs. CRF-GNNs inserts CRF layer between the graph convolutional layers, which regularizes GNNs to preserve the label dependencies among nodes (Gao et al., 2019). SPN (Qu et al., 2022) models the local label correlation of each linked node pair via nodeGNN and edgeGNN and takes the edges in graph as input to propagate all the pairwise label correlations along edges. Difference compared to existing work: ① While CRF-based methods focus on modeling the local label correlation of every linked node pair, we learn the global joint distribution of node and its cluster. ② Our proposals show promising training scalability and inference efficiency. The CRF-based methods take the whole graph as input to propagate all the pairwise label correlations along edges. In contrast, we train and infer the joint distribution of target node only with one reference signal (i.e., cluster), which allows the batch training on large graphs (e.g., Amazon with millions of nodes).
@@ 5 Experiments
In this section, we evaluate our joint-cluster learning framework on 12 public datasets over 7 backbone models. The code is available at: https://github.com/MR9812/Joint-Cluster-Supervised-Learning
@@ Appendix A Algorithm
The detailed description of our proposed joint-cluster learning framework.